OpenStudy (anonymous):

I need help figuring out this problem! Write an expression for the function f(x) if it is known that f'(x)=cos(x^2) and f(1)=7 f(x)=C+integral from a to x A(t)dt C= a= A(t)= cos(t^2)

OpenStudy (anonymous):

\[f(x)=C+\int\limits_{a}^{x} A(t)dt\]

ganeshie8 (ganeshie8):

use FTC

OpenStudy (abb0t):

@Squirrels

ganeshie8 (ganeshie8):

\[\large f(x) - f(a) = \int \limits_a^x f'(t) dr\]

ganeshie8 (ganeshie8):

plugin \(a = 1\) and the given function \(f'(x) \)

ganeshie8 (ganeshie8):

\[\large f(x) - f(1) = \int \limits_1^x \cos (t^2) dt\] \[\large f(x) = f(1) + \int \limits_1^x \cos (t^2) dt\] \[\large f(x) = 7 + \int \limits_1^x \cos (t^2) dt\]